OpenStudy (anonymous):

What is the abscissa of the midpoint of the line segment whose endpoints are

OpenStudy (anonymous):

\[(5\sqrt{2},2\sqrt{3}) \space and \space (\sqrt{2},2\sqrt{3}) \space ?\]

OpenStudy (anonymous):

is it \[2\sqrt{3}\]

OpenStudy (lgbasallote):

abscissa is the x-coodinate

OpenStudy (lgbasallote):

...so in short it's impossible to be that

OpenStudy (lgbasallote):

because the abscissa of the midpoint should be \[\huge \frac{x_1 +x_2}{2}\]

OpenStudy (lgbasallote):

\[\huge \implies \frac{5\sqrt 2 + \sqrt 2}{2}\] make sense?

OpenStudy (anonymous):

3\[\sqrt{2}\]?

OpenStudy (lgbasallote):

yes